OpenStudy (anonymous):

Write the standard equation of a circle with each given radius and center. r = 4; C= (0, 0) How do you do this?

OpenStudy (anonymous):

Hi Shantia, First, the equation of a circle is (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

(h,k) is known as the center, as it is the vertical and horizontal shift.

OpenStudy (zehanz):

Standard equation of a circle with center (a, b) and radius r is: (x-a)²+(y-b)²=r². Here, you have (a, b)=(0, 0) and r = 4.

OpenStudy (anonymous):

here the center is (0,0), thus h =0 and k=0. R = 4. Thus your equation is (x-0)^2+ (y-0)^2 = 4^2

OpenStudy (anonymous):

Thanks guys!

OpenStudy (zehanz):

@Mathexpert: Good explaining, but it is better to leave something do do for the asker...

OpenStudy (anonymous):

No problem. Good luck.

OpenStudy (anonymous):

Do I do the same for this one "r = 1; C = (2, 3)" or is it different because they are not zeros?

OpenStudy (zehanz):

Now, (a, b) = (2, 3), so set a=2 and b=3 (and r=1) in the general equation!

OpenStudy (anonymous):

ok

OpenStudy (zehanz):

What equation do you get?

OpenStudy (anonymous):

(x-2)^2+(y-3)^2=1 ???

OpenStudy (zehanz):

Well done! You're getting it!

OpenStudy (anonymous):

Write the standard equation for the circle given the following equations. x2 + y2 - 8x + 7 = 0 im stuck again

OpenStudy (zehanz):

Complete the square: x² + y² - 8x + 7 = 0 x² - 8x + y² +7 = 0 (x-4)² - 16 + y² +7 = 0 (x-4)² + y² = 9

OpenStudy (anonymous):

I'm lost what happened to the seven and the sixteen?

OpenStudy (zehanz):

-16+7=-9. That is on the left of the equal sign. So add 9 to both sides to get ... = 9

OpenStudy (anonymous):

ahhhhh ok I see now thanks

OpenStudy (zehanz):

YW!

OpenStudy (anonymous):

x2 + y2 + 4x - 6y - 3 = 0 so would this be my answer (x-2)2 + (y-3)2 = 3

OpenStudy (anonymous):

(x-2^)2 + (y-3)^2 = 3

OpenStudy (zehanz):

OK, let me check: x² + y² + 4x - 6y - 3 = 0 x² + 4x + y² - 6y - 3 = 0 (x + 2)² - 4 + (y - 3)² - 9 -3 = 0 (x + 2)² + (y - 3)² - 4 - 9 -3 = 0 (x + 2)² + (y - 3)² - 16 = 0 (x + 2)² + (y - 3)² = 16 You have to complete the squares and then take all the constan terms together and put them on the other side of the equal sign.

OpenStudy (zehanz):

If you look well, you can see I replace: x² + 4x with (x + 2)² - 4 and y² - 6y with (y - 3)² - 9. This is because (x + 2)² = x² +4x + 4, and we only need the first two terms, not the "+4". So we have to subtract 4.